1/AB = 1/BC - 1/CD solve for C
Found 2 solutions by ikleyn, Alan3354:
Answer by ikleyn(52781)

.
Just solved at this link
https://www.algebra.com/algebra/homework/Expressions-with-variables/Expressions-with-variables.faq.question.1105881.html

https://www.algebra.com/algebra/homework/Expressions-with-variables/Expressions-with-variables.faq.question.1105881.html


Answer by Alan3354(69443)

1/AB = 1/BC - 1/CD solve for C
=================
1/AB = (1/C)*(1/B - 1/D) = (1/C)*((D-B)/BD)
C/AB = (D-B)/BD
C/A = (D-B)/D
C = (AD-AB)/D